Norcal issued five-year 7% bonds with a face value of $100,000, for $96,567.94 on January 1, Year 1 when the market effective rate of interest was 7.5%. The bonds pay annual interest each December 31. Stanton uses the effective interest method for amortization of premium or discount on bonds payable.

Required:

a) What is the annual amount of cash that Stanton will pay to bondholders for interest?

b) What amount of interest expense and discount amortization should Stanton recognize for Year 1? What is the carrying amount of the liability on December 31, Year 1?

c) What amount of interest expense and premium amortization should Stanton recognize for Year 2? What is the carrying amount of the liability on December 31, Year 2?

d) What is the total amount of interest that Stanton will record in interest expense over the life of the bond?
Answer #1

a) Annual amount of cash paid = 100000*7% = 7000

b) Interest expense = 96567.94*7.5% = 7242.60

Discount amortization = 7242.60-7000 = 242.60

Carrying amount = 96567.94+242.60 = 96810.54

c) Interest expense = 96810.54*7.5% = 7260.79

Discount amortization = 7260.79-7000 = 260.79

Carrying amount = 96810.54+260.79 = 97071.33

d) Total interest expense = 7000*5+(100000-96567.94) = 38432.06